# Generalized $T\bar{T}$-like flows for scalar theories in two dimensions

### v1

@@
H. Babaei-Aghbolagh,1,2,3 Song He,1,2,4,5 and Hao Ouyang4
@@
3Department of Physics, University of Mohaghegh Ardabili, P.O. Box 179, Ardabil, Iran
4Center for Theoretical Physics and College of Physics, Jilin University, Changchun 130012, China
5Max Planck Institute for Gravitational Physics (Albert Einstein Institute),
@@
E-mails: h.babaei@uma.ac.ir, hesong@nbu.edu.cn, haoouyang@jlu.edu.cn
We demonstrate that the necessary condition for $SO(N)\times SO(N)$ duality invariance manifests as a partial differential equation in two-dimensional scalar theories. This condition, expressed as a partial differential equation, corresponds precisely to the integrability condition. We derive a general perturbation solution to this partial differential equation, which includes both a root $T\bar{T}$ flow equation and an irrelevant $T\bar{T}$-like flow equation. Additionally, we identify a general form for these flow equations that commute with each other.
@@ 1 Introduction
In [8], a two-dimensional ModMax theory was introduced, which presents the root-type $T\bar{T}$ deformation in this context. This deformation was independently identified in [9, 10]. Studies highlight connections between $T\bar{T}$-like flows and gravity theories in various dimensions, particularly in models coupled to gravity, using the vielbein formalism for deformed field theories [11, 12, 13, 14, 15]. For recent advances on root-type $T\bar{T}$ deformation, see [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Ref. [9] demonstrates that operators (1.3) and (1.5) commute, as illustrated in Fig. 1.
@@
The solutions to this equation are theories exhibiting SO(2) symmetry. For instance, the nonlinear electrodynamic theories of Born-Infeld [33, 34], ModMax [7], and General ModMax [35, 36] are all solutions to the PDE 1.9. The relationship between electromagnetic duality and $T\bar{T}$-like deformations at the effective action level has been explored in [37, 38, 39, 40, 41, 42, 43, 44, 45]. Additionally, innovative methods for solving this equation can be found in [46, 47, 48, 49, 50]. The approach has also been extended to the Carrollian framework [51], where a Carrollian electrodynamics theory with both conformal invariance and SO(2) duality is explored by solving a PDE. In [52, 53], sigma models were introduced that are invariant under U(1) ”duality rotations,” which exchange the dynamical variables and their equations of motion. The Lagrangians of these sigma models obey a partial differential equation analogous to the self-duality equation followed by U(1) duality invariant models in nonlinear electrodynamics.
@@
T-duality, a central concept in perturbative string theory, emerges when the theory is compactified on a torus [54]. This compactification reveals that the spectrum of the free string on a torus remains invariant under $O(N,N)$ transformations [55, 56]. This symmetry extends beyond the free string spectrum to the full bosonic string theory, where compactification on a torus $T^{N}$ preserves invariance under $O(N,N)$ transformations [57]. More notably, after integrating the massive modes, T-duality manifests as symmetry in the effective actions [58]. It has been demonstrated that the dimensional reduction of the classical effective actions of bosonic string theory at each order in $\alpha^{\prime}$ is invariant under $O(N,N)$ transformations [58, 59]. This insight has led to the development of sigma models, where the worldsheet string theory remains invariant under $O(N,N)$ transformations in two dimensions. When the target space metric is $G_{ij}=\delta_{ij}$, and the Kalb-Ramond field vanishes, the symmetry group $O(N,N)$ reduces to $SO(N)\times SO(N)$. In Section 2.2, we show that the necessary and sufficient condition for invariance under $SO(N)\times SO(N)$ transformations is given by the following PDE, which involves the two Lorentz-invariant variables $P_{1}$ and $P_{2}$:
$0\displaystyle 8(P_{1}^{2}-P_{2})\left(\frac{\partial\mathcal{L}}{\partial P_{2}}\right)^{2}+8P_{1}\frac{\partial\mathcal{L}}{\partial P_{2}}\frac{\partial\mathcal{L}}{\partial P_{1}}+4\left(\frac{\partial\mathcal{L}}{\partial P_{1}}\right)^{2}-1=$ 0 . (1.10)
This PDE provides a unified solution framework for both free and deformed theories, such as 1.1, 1.4, 1.6, and 1.7 (see Refs. [60, 61]), which are classified together due to their adherence to the condition in 1.10.
Refs. [60, 61] demonstrate that the model achieves classical integrability when the Lagrangian satisfies Eq. 1.10. Therefore, the criterion for classical integrability of the model, as outlined in condition 1.10, coincides with the PDE that a Lagrangian must obey in a four-dimensional theory of duality-invariant electrodynamics.
@@
This paper aims to identify two categories of general solutions to the PDE 1.10, corresponding to distinct marginal and irrelevant $T\bar{T}$-like deformations. We will delineate the general structure of these deformations using marginal and irrelevant flow equations. Specifically, we classify theories of non-linear sigma models dependent on two Lorentzian variables, $P_{1}$ and $P_{2}$, which feature two types of coupling parameters: a dimensionless coupling $\gamma$ and a dimensional coupling $\lambda$. We will derive expressions for the $\gamma$ and $\lambda$ couplings from the root-type and irrelevant transformations of $T\bar{T}$.
@@
The organization of this paper is as follows: In Section 2, we show that the invariance condition under $SO(N)\times SO(N)$ is governed by the PDE in 1.10, which is identical to the equation derived from the integrability condition in Equation 3.32 of Ref. [60]. In Section 3, we solve the PDE in 1.10 using perturbation theory. This method considers all possible solutions in integer powers of $P_{1}$ and $P_{2}$ at each order of $\lambda$, determining the associated constants. We demonstrate that an irrelevant flow equation exists for this solution, constructed from integer powers of the $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$ structures, and is commutative with root deformation. In Section 4, we transform the PDE 1.10 into a new differential form using the variables $P_{1}$ and $\mathcal{P}$, where $\mathcal{P}:=\sqrt{-P_{1}^{2}+2P_{2}}$. This transformation allows us to derive a general perturbation solution and to determine how the solution coefficients depend on the $\gamma$ coupling, yielding a root flow equation. In Section 5, we identify the general form of the irrelevant flow equations for these theories, which commute with the root deformation. Finally, in Section 6, we summarize our results and discuss future directions.
@@
The integrability condition in Eq. 1.10 has recently been identified as a necessary and sufficient condition for $O(N,N)$ duality in two-dimensional scalar theories. These theories play a central role in integrable systems and are intimately connected with the integrable models discussed in this section. We explore how the invariance under $SO(N)\times SO(N)$ transformations emerges as a crucial feature for maintaining integrability and investigate the relationship between these two concepts in the context of the scalar theory.
@@ 2.1 Integrability condition as a PDE
Refs. [60, 61] present a detailed stress tensor analysis within a generic interacting chiral boson theory. The study carefully examines the classical flow properties governed by the stress tensor $T_{\mu\nu}$, focusing on maintaining Lorentz invariance. In Ref. [61], the authors derive a Lorentz-invariant condition for the Lagrangian of the principal chiral model (PCM).  [^1] The Lorentz condition for chiral scalar theory is given in Eq. (2.35) of Ref. [61] as $\bigl{(}\frac{\partial\mathcal{L}}{\partial S}\bigr{)}^{2}+2\frac{S}{P}\frac{\partial\mathcal{L}}{\partial S}\frac{\partial\mathcal{L}}{\partial P}+\bigl{(}\frac{\partial\mathcal{L}}{\partial P}\bigr{)}^{2}-1=0$. This is reformulated in Eq. 1.10 by substituting $S\Longleftrightarrow-\frac{1}{2}P_{1}$ and $P^{2}\Longleftrightarrow\frac{P_{1}^{2}-P_{2}}{2}$. The variables $S$ and $P$ are defined as $S=\frac{1}{2}\left(\phi^{\prime j}\phi^{\prime j}+\bar{\phi}^{\prime\bar{j}}\bar{\phi}^{\prime\bar{j}}\right)$ and $P=\frac{1}{2}\left(\phi^{\prime j}\phi^{\prime j}-\bar{\phi}^{\prime\bar{j}}\bar{\phi}^{\prime\bar{j}}\right)$. This condition is crucial for ensuring that the interaction function accurately reflects a Lorentz-invariant theory.
@@
$j=g^{-1}dg,\quad\quad\quad\tilde{j}=-(dg)g^{-1}\,,$ (2.1)
@@
Any Lagrangian that depends on $j_{\pm}$ only through the combinations $\text{tr}[j_{+}j_{-}]$ and $\text{tr}(j_{+}j_{+})\text{tr}[j_{-}j_{-}]$ can be expressed, after a change of variables, as a function $\mathcal{L}(P_{1},P_{2})$ of the two variables:
$P_{1}:=-\text{tr}[j_{+}j_{-}]\,,\qquad P_{2}:=\frac{1}{2}\left(\text{tr}[j_{+}j_{+}]\text{tr}[j_{-}j_{-}]+\left(\text{tr}[j_{+}j_{-}]\right)^{2}\right)\,,$ (2.5)
@@
$\mathfrak{L}_{\pm}^{(\lambda,\gamma)}=\frac{j_{\pm}\pm z\,\mathfrak{J}_{\pm}}{1-z^{2}}\,,$ (2.7)
@@
The flatness condition in equation 2.8 is met when the current $\mathfrak{J}_{\alpha}$ satisfies the property $\big{[}\mathfrak{J}_{+},j_{-}\big{]}=\big{[}j_{+},\mathfrak{J}_{-}\big{]}$.As shown in Ref. [60], this equality holds if and only if the Lagrangian $\mathcal{L}(P_{1},P_{2})$ is satisfied the PDE in 1.10.
@@ 2.2 $SO(N)\times SO(N)$ duality invariance
It is well known the worldsheet string theory remains invariant under $O(N,N)$ transformations in two dimensions. This invariance has led to the development of sigma models [55, 56]. Two-dimensional non-linear sigma models admit $O(N,N)$ duality transformations (a review can be found in [54]), which includes T-duality as a special case. When the target space metric $G_{ij}=\delta_{ij}$ and the Kalb-Ramond field is zero, a non-linear sigma model is invariant under a $SO(N)\times SO(N)$ subgroup duality transformation. In this subsection, we show that the condition 1.10 leads to the $SO(N)\times SO(N)$ duality invariance for scalar theories ${\cal L}(P_{1},P_{2})$ with $G_{ij}=\delta_{ij}$. We denote
@@
where the convention of Levi-Civita symbol is $\epsilon_{01}=1$. In the following, we will use matrix notation. For instance, we denote $\eta^{\mu\nu}\partial_{\mu}\phi^{i}\partial_{\nu}\phi^{i}=\mathrm{Tr}(F^{t}\eta^{-1}F)$ and 2.10 can be written as
@@
where $O_{1}$ and $O_{2}$ are two $N\times N$ special orthogonal matrices. The diagonal subgroup defined by $O_{1}=O_{2}=O$ is the usual $SO(N)$ symmetry $\phi^{j}\rightarrow\phi^{i}O_{i}^{~{}j}$. The genuine duality can be obtained as the coset modulo the diagonal subgroup. When $N=2$, this duality transformation is related to the duality of four-dimensional electrodynamics via dimensional reduction. In the treatment of duality, $F$ and $H$ are not seen as functionals of $\phi$, and are constrained by 2.10 or equivalently 2.11, where $\mathcal{L}$ is regarded a function of $F$.
The duality invariance requires that 2.11 is invariant under the transformations 2.14. We now show that 1.10 leads to duality invariance. For an arbitrary $\mathcal{L}$, $H$ can be computed as
@@
To derive the formula, we use the observation that $FbF^{t}$ is a $2\times 2$ antisymmetric matrix, and therefore, the trace can be factorized. When the condition 1.10 is satisfied, we find
@@
Therefore, the constraint 2.11 is invariant under the duality rotation, and the condition 1.10 leads to the $SO(N)\times SO(N)$ duality invariance.
In [62], it was proved that the T-duality transformation commutes with stress-energy tensor transformations using an auxiliary field formulation [13, 63]. This result can be easily extended to the $O(N,N)$ duality transformations. Consequently, stress-energy tensor transformations preserve the $SO(N)\times SO(N)$ duality invariance, which explains that stress-energy tensor transformations can realize perturbative solutions of 1.10.

### v2

@@
H. Babaei-Aghbolagh,1,2 Song He,1,2,3,4 and Hao Ouyang3
@@
3Center for Theoretical Physics and College of Physics, Jilin University, Changchun 130012, China
4Max Planck Institute for Gravitational Physics (Albert Einstein Institute),
@@
E-mails: hosseinbabaei@nbu.edu.cn, hesong@nbu.edu.cn, haoouyang@jlu.edu.cn
We demonstrate that the necessary condition for $SO(N)\times SO(N)$ duality invariance manifests as a partial differential equation in two-dimensional scalar theories. This condition, expressed as a partial differential equation, corresponds precisely to the integrability condition. We derive a general perturbation solution to this partial differential equation, which includes both a root $T\bar{T}$ flow equation and an irrelevant $T\bar{T}$-like flow equation. Additionally, we identify a general form for these flow equations that commute with each other. Our results establish a general integrable theory characterized by theory-dependent coefficients at each order in the $\lambda$-expansion. This unified framework systematically classifies all integrable theories possessing two Lorentz-invariant variables ($P_{1}$, $P_{2}$) while accommodating arbitrary orders of the coupling constants ($\lambda$, $\gamma$). The theory provides a comprehensive classification scheme that encompasses both known and novel integrable systems within this class.
@@ 1 Introduction
After presenting the ModMax theory in four dimensions in Ref. [7] and the flow equation based on the $\gamma$ coupling in Ref. [6], the initial version of Ref. [8] introduced a two-dimensional ModMax theory. Subsequently, three papers appeared in arXiv : the second version of Ref. [8] and Ref. [9] were appeared on the same day, followed five days later by Ref. [10], which presented the $\gamma$-flow equation for the two-dimensional ModMax theory in the form of operator (1.5). Studies highlight connections between $T\bar{T}$-like flows and gravity theories in various dimensions, particularly in models coupled to gravity, using the vielbein formalism for deformed field theories [11, 12, 13, 14, 15]. For recent advances on root-type $T\bar{T}$ deformation, see [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@
Refs. [10, 9] demonstrates that operators (1.3) and (1.5) commute, as illustrated in Fig. (1).
@@
The solutions to this equation are theories exhibiting SO(2) symmetry. For instance, the nonlinear electrodynamic theories of Born-Infeld [35, 36], ModMax [7], and General ModMax [37, 38] are all solutions to the PDE (1.9). The relationship between electromagnetic duality and $T\bar{T}$-like deformations at the effective action level has been explored in [39, 40, 41, 42, 43, 44, 45, 46, 47]. Additionally, innovative methods for solving this equation can be found in [48, 49, 50, 51, 52, 53, 54, 55]. The approach has also been extended to the Carrollian framework [56], where a Carrollian electrodynamics theory with both conformal invariance and SO(2) duality is explored by solving a PDE. In [57, 58], sigma models were introduced that are invariant under U(1) ”duality rotations,” which exchange the dynamical variables and their equations of motion. The Lagrangians of these sigma models obey a partial differential equation analogous to the self-duality equation followed by U(1) duality invariant models in nonlinear electrodynamics.
@@
T-duality, a central concept in perturbative string theory, emerges when the theory is compactified on a torus [59]. This compactification reveals that the spectrum of the free string on a torus remains invariant under $O(N,N)$ transformations [60, 61]. This symmetry extends beyond the free string spectrum to the full bosonic string theory, where compactification on a torus $T^{N}$ preserves invariance under $O(N,N)$ transformations [62]. More notably, after integrating the massive modes, T-duality manifests as symmetry in the effective actions [63]. It has been demonstrated that the dimensional reduction of the classical effective actions of bosonic string theory at each order in $\alpha^{\prime}$ is invariant under $O(N,N)$ transformations [63, 64]. This insight has led to the development of sigma models, where the worldsheet string theory remains invariant under $O(N,N)$ transformations in two dimensions. When the target space metric is $G_{ij}=\delta_{ij}$, and the Kalb-Ramond field vanishes, the symmetry group $O(N,N)$ reduces to $SO(N)\times SO(N)$. In Section (2.2), we show that the necessary and sufficient condition for invariance under $SO(N)\times SO(N)$ transformations is given by the following PDE, which involves the two Lorentz-invariant variables $P_{1}$ and $P_{2}$:
$8(P_{1}^{2}-P_{2})\left(\frac{\partial\mathcal{L}}{\partial P_{2}}\right)^{2}+8P_{1}\frac{\partial\mathcal{L}}{\partial P_{2}}\frac{\partial\mathcal{L}}{\partial P_{1}}+4\left(\frac{\partial\mathcal{L}}{\partial P_{1}}\right)^{2}-1=0.$ (1.10)
This PDE provides a unified solution framework for both free and deformed theories, such as (1.1), (1.4), (1.6), and (1.7) (see Refs. [65, 66]), which are classified together due to their adherence to the condition in (1.10). Refs. [65, 66] demonstrate that the model achieves classical integrability when the Lagrangian satisfies Eq. (1.10). Therefore, the criterion for classical integrability of the model, as outlined in condition (1.10), coincides with the PDE that a Lagrangian must obey in a four-dimensional theory of duality-invariant electrodynamics. The integrability condition for all integrable theories composed of two variables $P_{1}$ and $P_{2}$ manifests as the differential equation presented in Eq. (1.10) to all orders in $\lambda$. Notably, both ModMax theory and the generalized Born-Infeld theory emerge as particular cases within this broader class of theories. In this work, we introduce a novel systematic approach for generating all such integrable theories by solving the integrability condition given in Eq. (1.10). Our methodology possesses the following key features:
  • In the absence of computational constraints, this approach yields exact, integrable solutions to all orders in $\lambda$ through a deformation procedure
  • The perturbative solution can provide a framework for understanding integrable theories in this class.
  • We explicitly derive a family of non-trivial solutions through $\mathcal{O}(\lambda^{n})$, where $n$ is an arbitrary integer
@@
This paper aims to identify two categories of general solutions to the PDE (1.10), corresponding to distinct marginal and irrelevant $T\bar{T}$-like deformations. We will delineate the general structure of these deformations using marginal and irrelevant flow equations. Specifically, we classify theories of non-linear sigma models dependent on two Lorentzian variables, $P_{1}$ and $P_{2}$, which feature two types of coupling parameters: a dimensionless coupling $\gamma$ and a dimensional coupling $\lambda$. We will derive expressions for the $\gamma$ and $\lambda$ couplings from the root-type and irrelevant transformations of $T\bar{T}$.
@@
The organization of this paper is as follows: In Section (2), we show that the invariance condition under $SO(N)\times SO(N)$ is governed by the PDE in (1.10), which is identical to the equation derived from the integrability condition in Equation 3.32 of Ref. [65]. In Section (3), we solve the PDE in (1.10) using perturbation theory. This method considers all possible solutions in integer powers of $P_{1}$ and $P_{2}$ at each order of $\lambda$, determining the associated constants. We demonstrate that an irrelevant flow equation exists for this solution, constructed from integer powers of the $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$ structures, and is commutative with root deformation. In Section (4), we transform the PDE (1.10) into a new differential form using the variables $P_{1}$ and $\mathcal{P}$, where $\mathcal{P}:=\sqrt{-P_{1}^{2}+2P_{2}}$. This transformation allows us to derive a general perturbation solution and to determine how the solution coefficients depend on the $\gamma$ coupling, yielding a root flow equation. In Section (5), we identify the general form of the irrelevant flow equations for these theories, which commute with the root deformation. Finally, in Section (6), we summarize our results and discuss future directions.
@@
The integrability condition in Eq. (1.10) has recently been identified as a necessary and sufficient condition for $O(N,N)$ duality in two-dimensional scalar theories. These theories play a central role in integrable systems and are intimately connected with the integrable models discussed in this section. We explore how the invariance under $SO(N)\times SO(N)$ transformations emerges as a crucial feature for maintaining integrability and investigate the relationship between these two concepts in the context of the scalar theory.
@@ 2.1 Integrability condition as a PDE
Refs. [65, 66] present a detailed stress tensor analysis within a generic interacting chiral boson theory. The study carefully examines the classical flow properties governed by the stress tensor $T_{\mu\nu}$, focusing on maintaining Lorentz invariance. In Ref. [66], the authors derive a Lorentz-invariant condition for the Lagrangian of the principal chiral model (PCM).  [^1] The Lorentz condition for chiral scalar theory is given in Eq. (2.35) of Ref. [66] as $\bigl{(}\frac{\partial\mathcal{L}}{\partial S}\bigr{)}^{2}+2\frac{S}{P}\frac{\partial\mathcal{L}}{\partial S}\frac{\partial\mathcal{L}}{\partial P}+\bigl{(}\frac{\partial\mathcal{L}}{\partial P}\bigr{)}^{2}-1=0$. This is reformulated in Eq. (1.10) by substituting $S\Longleftrightarrow-\frac{1}{2}P_{1}$ and $P^{2}\Longleftrightarrow\frac{P_{1}^{2}-P_{2}}{2}$. The variables $S$ and $P$ are defined as $S=\frac{1}{2}\left(\phi^{\prime j}\phi^{\prime j}+\bar{\phi}^{\prime\bar{j}}\bar{\phi}^{\prime\bar{j}}\right)$ and $P=\frac{1}{2}\left(\phi^{\prime j}\phi^{\prime j}-\bar{\phi}^{\prime\bar{j}}\bar{\phi}^{\prime\bar{j}}\right)$. This condition is crucial for ensuring that the interaction function accurately reflects a Lorentz-invariant theory.
@@
$j=g^{-1}dg,\quad\quad\quad\tilde{j}=-(dg)g^{-1}\,.$ (2.1)
@@
Any Lagrangian that depends on $j_{\pm}$ only through the combinations of $\text{tr}[j_{+}j_{-}]\,\,$ and $\text{tr}(j_{+}j_{+})\,\,\text{tr}[j_{-}j_{-}]$ can be expressed, after a change of variables, as a function $\mathcal{L}(P_{1},P_{2})$ of the two variables:
$P_{1}:=-\text{tr}[j_{+}j_{-}]\,,\qquad P_{2}:=\frac{1}{2}\left(\text{tr}[j_{+}j_{+}]\text{tr}[j_{-}j_{-}]+\left(\text{tr}[j_{+}j_{-}]\right)^{2}\right)\,.$ (2.5)
@@
$\mathfrak{L}_{\pm}^{(\lambda,\gamma)}=\frac{j_{\pm}\pm z\,\mathfrak{J}_{\pm}}{1-z^{2}}\,.$ (2.7)
@@
The flatness condition in equation (2.8) is met when the current $\mathfrak{J}_{\alpha}$ satisfies the property $\big{[}\mathfrak{J}_{+},j_{-}\big{]}=\big{[}j_{+},\mathfrak{J}_{-}\big{]}$.As shown in Ref. [65], this equality holds if and only if the Lagrangian $\mathcal{L}(P_{1},P_{2})$ is satisfied the PDE in (1.10).
@@ 2.2 $SO(N)\times SO(N)$ duality invariance
It is well known the worldsheet string theory remains invariant under $O(N,N)$ transformations in two dimensions. This invariance has led to the development of sigma models [60, 61]. Two-dimensional non-linear sigma models admit $O(N,N)$ duality transformations (a review can be found in [59]), which includes T-duality as a special case. When the target space metric $G_{ij}=\delta_{ij}$ and the Kalb-Ramond field is zero, a non-linear sigma model is invariant under a $SO(N)\times SO(N)$ subgroup duality transformation. In this subsection, we show that the condition (1.10) leads to the $SO(N)\times SO(N)$ duality invariance for scalar theories ${\cal L}(P_{1},P_{2})$ with $G_{ij}=\delta_{ij}$. We denote
@@
where the convention of Levi-Civita symbol is $\epsilon_{01}=1$. In the following, we will use matrix notation. For instance, we denote $\eta^{\mu\nu}\partial_{\mu}\phi^{i}\partial_{\nu}\phi^{i}=\mathrm{Tr}(F^{t}\eta^{-1}F)$ and (2.10) can be written as
@@
where $O_{1}$ and $O_{2}$ are two $N\times N$ special orthogonal matrices. The diagonal subgroup defined by $O_{1}=O_{2}=O$ is the usual $SO(N)$ symmetry $\phi^{j}\rightarrow\phi^{i}O_{i}^{~j}$. The genuine duality can be obtained as the coset modulo the diagonal subgroup. When $N=2$, this duality transformation is related to the duality of four-dimensional electrodynamics via dimensional reduction. In the treatment of duality, $F$ and $H$ are not seen as functionals of $\phi$, and are constrained by (2.10) or equivalently (2.11), where $\mathcal{L}$ is regarded a function of $F$.
The duality invariance requires that (2.11) is invariant under the transformations (2.14). We now show that (1.10) leads to duality invariance. For an arbitrary $\mathcal{L}$, $H$ can be computed as
@@
To derive the formula, we use the observation that $FbF^{t}$ is a $2\times 2$ antisymmetric matrix, and therefore, the trace can be factorized. When the condition (1.10) is satisfied, we find
@@
Therefore, the constraint (2.11) is invariant under the duality rotation, and the condition (1.10) leads to the $SO(N)\times SO(N)$ duality invariance.
In [67], it was proved that the T-duality transformation commutes with stress-energy tensor transformations using an auxiliary field formulation [13, 68]. This result can be easily extended to the $O(N,N)$ duality transformations. Consequently, stress-energy tensor transformations preserve the $SO(N)\times SO(N)$ duality invariance, which explains that stress-energy tensor transformations can realize perturbative solutions of (1.10).